Ticket Number: HEA-NOC/20040118-2 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20040118 16:21 UTC Problem Start: 20040118 15:14 UTC
Ticket Closed: Problem End:
Site/line: MLE (HEAN0108)
1 minute outage on MLE circuit.
All traffic to and from MLE.
20040118 16:13 JL
Called Esat to report fault, they saw similiar errors on circuit at 15:14.
MLE called Esat to report that theyre working on their router. OSPF
working over primary and backup circuit however secondary interface IP no
longer reponsing to pings. Will continue to monitor.
20040118 19:31 JL
A number of further OSPF adjacency changes, primary circuit now in init
state, secondary in full. Traffic flowing using backup circuit. Will
require further investigation and intervention on 20040119 to have primary
circuit brought up.
This ticket can be monitored at http://www.hea.net/tickets/20040118-2